OpenStudy (anonymous):

you know that 2x5 = 5x2. is it true that 2 to the power of 5 = 5 to the power of 2? explain. If this is not true, which is greater?

OpenStudy (anonymous):

compute and see what you get \[2^5=2\times 2\times 2\times 2\times 2\] whereas \[5^2=5\times 5\]

OpenStudy (anonymous):

The fundamental theorem of arithmetic states that they can't be equal because their prime factorization is not the same.

OpenStudy (anonymous):

Basically exponentiation is not commutatative like multiplication. Different operations have different properties so what works for addition may not work for other operations.